A jet plane is flying at a certain altitude. At time t1 = 0, it has components of velocity Vx = 90 m/s, Vy = 110 m/s. At time t2 = 30 sec, the components are Vx = -170 m/s, Vy = 40 m/s. Calculate (a) the components of average acceleration, (b) the magnitude and direction of average acceleration.


1
Expert's answer
2021-07-12T12:11:59-0400

a)


"a_x=\frac{-170-90}{30}=-8.67\frac{m}{s^2}\\
a_y=\frac{40-110}{30}=-2.33\frac{m}{s^2}\\"

b)


"a=\sqrt{2.33^2+8.67^2}=9.0\frac{m}{s^2}\\\theta=180+\arctan{\frac{2.33}{8.67}}=195\degree"